OpenStudy (anonymous):

Help and explain please Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 -2x+2x=-6+6 0=0 probe :)

OpenStudy (anonymous):

2 3 All real numbers No solution which one?

OpenStudy (anonymous):

0x=0? eternal solutions :)

OpenStudy (anonymous):

is a indetermination!

OpenStudy (anonymous):

Lol so no solution?